500 reduced by 3 times my age is 278?
Answer by Theo(13342)

x = your age.
500 - 3x = 278
add 3x to both sides of this equation and subtract 278 from both sides of this equation to get:
500 - 278 = 3x
simplify to get 222 = 3x
solve for x to get x = 222 / 3 = 74
your age is 74.